# Stigmergic Swarming Agents for Fast Subgraph Isomorphism

### v1

@@ Abstract.
Swarming agents, Stigmergy, Subgraph Isomorphism, Ant colony optimization, ACO
@@ 1. Introduction
For example, chemists frequently seek shared structures in large organic molecules, where nodes are the atoms in a molecule (e.g., C, H, O, …), and edges are chemical bonds between them. Finding a common structure of six atoms between two organic molecules of 50 atoms each with a naïve enumeration requires over $10^{17}$ comparisons, motivating chemists to pursue research in subgraph isomorphism algorithms RN1.
@@ 2. Related Work
Population Single SolutionRN15; RN16; RN17; RN18; RN20; RN24; RN23; RN25; RN26 Robustness, Accessibility
@@ 4. The ASSIST Heuristic
ASSIST is inspired by the ACO TSP algorithm. Instead of Hamiltonian circuits, we seek paths between the query and data graphs that traverse matching fragments of both graphs (typically, a single edge). Figure 2 illustrates the movements of a single agent. Starting at a node labeled $A$ in the query, it seeks an $A$ node in the data that has similar neighbors to its original node (1). Then it looks for a neighbor in the data that matches a neighbor of $A$ in the query (B), returns to a $B$ node in the query, and seeks to complete the circuit. If it is successful, it has found a candidate edge in the desired matching, and augments the pheromones on the nodes and edges in both the query and the data.
@@
In outlining ASSIST, we discuss
  • the data model for the graphs we analyze,
  • how the graphs are initialized when loaded into ASSIST,
  • the peering process that identifies possible matches of individual nodes between the query and the data graphs, and
  • the process of matching a single edge.
@@ 4.2. Ontology and Pseudocode
ASSIST is implemented in Repast North2013, in which the basic unit of time is the tick. Algorithm 1 shows the sequence of actions in each tick.
@@ 4.3. Initialization
Next, $pherLevel\leftarrow 1.0$ on all retained nodes, allowing them initialize $nbrPhers$. After agents update pheromones on visited nodes (Algorithm 2 line 29, invoked in Algorithm 1 line 9), nodes update $nbrPher$s (Algorithm 1, line 11).
@@ 4.5. Detecting Termination and Retrieving Results
ASSIST’s matching process is continuous and emergent. For it to be useful in practice, we need to do three things:
  1. (1)
    measure its performance on problems of varying complexity;
  2. (2)
    tell when it has converged;
  3. (3)
    retrieve the matching subgraph it has found.
@@
The next two figures show the same run, terminated when the number of matched edges has been stable for 20 ticks. The kernel is discovered at tick 8 (the vertical dashed line). The run discovers not only the 10-node, 16-edge kernel, but four more nodes and edges that are connected with it in the same way in the query and data.
@@ 5.1. Dimensions of Interest
  • Node Ambiguity: In our baseline experiments, within each graph, each node’s label-detail combination is unique, and matches require matching both label and detail. We expect matching time to increase as we increase the proportion of nodes in the query with detail = 0.
@@ 5.2. Data and Query Size
Figure 7 shows matching and peering time for a 100 node query as a function of data size. The inset (for data from 100 to 10k nodes) shows that from data size 2000 and up, matching time is basically flat, and the full graph shows that this holds for $10^{5}$ and $10^{6}$. Under these circumstances, peering dominates matching time for larger data.
Figure 8 shows kernel matching and peering times for a 4000 node data graph as a function of query size. The dependency in both cases is reasonably linear, though peering grows more slowly than matching. We expect this dependency. The larger the query, the larger the set of peered nodes and edges that the agents need to explore.
@@
We predicted peering time of $q\cdot log(d)$. Figure 8 shows the linear dependence on query size, but the variation in Figure 7 is too high for a useful fit of the $log(d)$ component. Figure 9 shows that peering time for equal-sized queries and data, of sizes {100, 1000, 2000, 3000, 4000, 5000, 6000, 7000, 8000, 9000, 10000} is indeed linear in $q\cdot log(d)$.
@@
5.3. Final Graph Size
@@ 5.4. Query Ambiguity
In many use cases, the query specifies the type of a node, but not its unique identity. Here we explore the effect of ignoring the details on some proportion of the query nodes, with a kernel of size 40, query of size 100, and data of size 6000. When we ignore the detail of a node, we say that we “ablate” it. Ablation requires the algorithm to consider more peers in the data graph for each node in the query than would be needed if we used the detail, and finds subgraphs that satisfy the category (label) of an ablated query node even if the query does not specify the detail.
@@
Figure 11 shows the impact of varying probabilities of ablation on two scenarios, one with label vocabulary 100 (the same size as other experiments reported here), the other with vocabulary of 10. In both cases, the matching time is linear, though with considerable variation, particularly for the smaller vocabulary. The matching time is much lower for the larger vocabulary. Because the labels even without detail carry ten times more information with the larger vocabulary, the peered data graph has fewer nodes and edges than with the smaller vocabulary, and presents an easier search problem.
@@ 6. Future Work
The experiments reported here show that ASSIST can find subgraphs in time $O(q\cdot log(d))$, but raise a number of questions that invite further study.
@@ 6.3. More Complex Matches
The query in Figure 2, repeated in Figure 12 (“Imprecise”), handles case 1 (here, matching ‘B’ with ‘b’) by allowing the agents to consult an ontology in case of mismatch to see if one of the nodes subsumes the other. The total pheromone deposited in the case of an imprecise match will be less than that deposited for an exact match.
@@ 7. Conclusion
ASSIST, a swarming stigmergic algorithm, offers an extremely rapid heuristic for subgraph isomorphism. After initial peering (which requires time $O(q\cdot log(d))$), matching is linear in query size and constant in data size, much faster than the best previous heuristics, which are quadratic in the number of nodes. In addition, it allows approximate matches, in which a query that specifies only a node’s category can retrieve subgraphs that match specific individuals in that category from the data.

### v2

@@ Abstract.
Swarming agents; Stigmergy; Subgraph Isomorphism; Ant colony optimization; ACO
doi:
DEZC1973
{CCSXML}
[…]
@@ 1. Introduction
For example, chemists frequently seek shared structures in large organic molecules, where nodes are the atoms in a molecule (e.g., C, H, O, …), and edges are chemical bonds between them. Finding a common structure of six atoms between two organic molecules, each of 50 atoms, with a naïve enumeration requires over $10^{17}$ comparisons, motivating chemists to pursue research in subgraph isomorphism algorithms RN1.
@@ 2. Related Work
Multiple Solutions Single Solution RN15; RN16; RN17; RN18; RN20; RN24; RN23; RN25; RN26 Robustness, Accessibility
@@ 4. The ASSIST Heuristic
Instead of Hamiltonian circuits, ASSIST seek paths between the query and data graphs that traverse matching fragments of both graphs (typically, a single edge). Figure 2 illustrates the movements of a single agent. Starting at a node labeled $A$ in the query, it seeks an $A$ node in the data that has similar neighbors to its original node (1). Then it looks for a neighbor in the data that matches a neighbor of $A$ in the query (B), returns to a $B$ node in the query, and seeks to complete the circuit. If it is successful, it has found a candidate edge in the desired matching, and augments the pheromones on the nodes and edges in both the query and the data.
@@
The following sections discuss 1) the data model for the graphs we analyze, 2) the classes and pseudocode of the algorithm, 3) how the graphs are initialized and possible matches identified, 4) the matching of a single edge, and 5) how we detect termination.
@@ 4.2. Classes and Pseudocode
ASSIST is implemented in Repast North2013, which measures time in ticks. Algorithm 1 shows the sequence of actions in each tick.
@@ 4.3. Initialization
Next, $pherLevel\leftarrow 1.0$ on all retained nodes, allowing them to initialize $nbrPhers$. Each time agents update pheromones on visited nodes (Algorithm 2 line 29, invoked in Algorithm 1 line 9), nodes update $nbrPher$s (Algorithm 1, line 11).
@@ 4.5. Detecting Termination and Retrieving Results
ASSIST’s matching process is continuous and emergent. For it to be useful in practice, we need to 1) measure its performance on problems of varying complexity, 2) tell when it has converged, and 3) retrieve the matching subgraphs it has found.
@@
Figures 4 and 6 show the same run, terminated when the number of matched edges has been stable for 20 ticks. The kernel is discovered at tick 8 (the vertical dashed line). The run discovers not only the 10-node, 16-edge kernel, but four more nodes and edges that are connected with it in the same way in the query and data.
@@ 5.1. Dimensions of Interest
  • Query Ambiguity: In our baseline experiments, within each graph, each node’s label-detail combination is unique, and matches require matching both label and detail. We expect matching time to increase as we increase the proportion of nodes in the query with detail = 0.
@@ 5.2. Data and Query Size
Figure 7 shows matching and peering time for a 100 node query as a function of data size d. The inset (for d from 100 to 10k nodes) shows that for $d>3000$, matching time is basically flat, and the full graph shows that this holds for $10^{5}$ and $10^{6}$. Under these circumstances, peering dominates matching time for larger data.
Figure 8 shows kernel matching and peering times for $d=4000$ as a function of query size q. The dependency in both cases is reasonably linear, though peering grows more slowly than matching. We expect this dependency. The larger $q$, the larger the set of peered nodes and edges that the agents need to explore.
@@
We predicted peering time of $q\cdot log(d)$. Figure 8 shows the linear dependence on $q$, but the variation in Figure 7 is too high for a useful fit of the $log(d)$ component. Figure 9 shows that peering time for $q=d\in\{100,1000,2000,3000,4000,5000,6000,7000,8000,9000,10000\}$ is indeed linear in $q\cdot log(d)$.
@@
5.3. Common Graph Size
@@ 5.4. Query Ambiguity
In many use cases, the query specifies the type of a node, but not its unique identity. Here we explore the effect of ignoring the details on some proportion of the query nodes, with a kernel of size 40, $q=100,d=6000$. When we ignore the detail of a node, we say that we “ablate” it. Ablation requires the algorithm to consider more peers in the data graph for each node in the query than would be needed if we used the detail, and finds subgraphs that satisfy the category (label) of an ablated query node even if the query does not specify the detail.
@@
Figure 11 shows the impact of varying probabilities of ablation on two scenarios, one with label vocabulary 100 (the same size as other experiments reported here), the other with vocabulary of 10. In both cases, matching time is linear, though variation is inverse to vocabulary. Because the labels even without detail carry ten times more information with the larger vocabulary, the peered data graph has fewer nodes and edges than with the smaller vocabulary, and presents an easier search problem.
@@ 6. Future Work
These experiments show that ASSIST can find subgraphs in time $O(q\cdot log(d))$, but leave room for further study.
6.1. Parallelization
The current code is single-threaded. The highly parallel actions of multiple agents hold great promise for multithreading and reimplementation on a GPGPU.
@@ 6.4. More Complex Matches
The query in Figure 2, repeated in Figure 12 (“Imprecise”), handles case 1 (here, matching ‘B’ with ‘b’) by having agents consult an ontology in case of mismatch to see if one of the nodes subsumes the other. The total pheromone deposited for an imprecise match will be less than that for an exact match.
@@ 7. Conclusion
ASSIST, a swarming stigmergic algorithm, offers an extremely fast heuristic for subgraph isomorphism. After initial peering (which requires time $O(q\cdot log(d))$), matching is linear in query size and constant in data size, much faster than the best previous heuristics, which are quadratic in the number of nodes. In addition, it allows approximate matches, in which a query that specifies only a node’s category can retrieve subgraphs that match specific individuals in that category from the data.